OpenStudy (anonymous):

Simplify completely 18st^4/52s^3t (times) 16s^3/9s^2t

OpenStudy (anonymous):

Is this your problem: \[\frac{18st^{4}}{52s^{3}t} \times \frac{16s^{3}}{9s^{2}t}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{18st^{4}}{52s^{3}t} \times \frac{16s^{3}}{9s^{2}t}\]\[\frac{18st^{4}}{52s^{3}t} \times \frac{16s^{3}}{9s^{2}t} = \frac{\cancel{18}2\cancel{s}t \cancel{^{4}}^{3}}{\cancel{52}13\cancel{s^{3}}t} \times \frac{\cancel{16}4\cancel{s^{3}}}{\cancel{9}s \cancel{^{2}}\cancel{t}} = \frac{2t \cancel{^{3}}^{2}}{13\cancel{t}} \times \frac{4}{s} = \frac{8t^{2}}{13s}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No Problem :)